OpenStudy (styxer):

Consider a parabola y = x² . The line that goes through the point (0,3/2) and is orthogonal to a tangent line to the part of parabola y=x² with x>0 is y= -x/2 + 3/2 , and x-coordinate of the intersection of the above two lines is [??]

OpenStudy (styxer):

So, to found the -1/2 I derivated the y=x² y'=2x and did m1 * m2 = -1 , applied in (y-3/2)= -1/2 * (x-0) to find y = -x/2 + 3/2 ... But calculating 2x = -x/2 + 3/2 , I don't find the correct answer of x-coordinate... why? :(

OpenStudy (stamp):



OpenStudy (loser66):

because the intersection point between y =x^2 and the line -x/2 + 3/2, not 2x with -x/2+3/2

OpenStudy (styxer):

@Loser66 ok but the question are saying "two lines", shouldn't we assume that these lines are what we just found?

OpenStudy (loser66):

common!! " line" is not used only for a straight one.

OpenStudy (styxer):

@Loser66 oh... doing x² = -x/2 + 3/2 I found the answer...

OpenStudy (loser66):

Tell me, what do you call the latitude line? a curve or a line?

OpenStudy (loser66):

ok, good

OpenStudy (styxer):

@Loser66 sorry, sometimes I get confused about these terms... English is not my native language ^_^'

OpenStudy (styxer):

thanks for your help

OpenStudy (loser66):

neither mine.:)